OpenStudy (anonymous):

At a high school, the average grade for the English 10 provincial is 64, with a standard deviation of 10. If 20 students scored between 73 and 86 on the exam, how many students took the exam? Use z scores in your solution

OpenStudy (kropot72):

Can you find the z-score for 73 marks using \[z=\frac{X-\mu}{\sigma}\]

OpenStudy (anonymous):

Excuse me?

OpenStudy (kropot72):

Do you know how to substitute the given values in \[z=\frac{X-\mu}{\sigma}\]

OpenStudy (anonymous):

I believe I do.

OpenStudy (kropot72):

So what z-scores do you get for marks of 73 and 86?

OpenStudy (anonymous):

\[z_{ 73}=\frac{73-64}{10}=0.9\] \[z_{ 86}=\frac{86-64}{10}=2.2\]

OpenStudy (kropot72):

So now it is nenecessary to find the percentage corresponding to the shaded area by referring to the standard normal distribution table with the values found for z  What percentages do you get for z = 0.9 and z = 2.2 ?